3 2
4 - - 2 -
5 3

Answer by stanbon!(97)

Convert the 4 3/5 to 5ths. "4" is 20/5. Add 3/5 to get 23/5
Convert the 2 2/3 to 3rds. "2" is 6/3. Add 2/3 to get 8/3
The problem is now the following:
23/5 - 8/3
Since the denominators are not the same determine the LCD of 5(3)= 15
Now convert each fraction to 15ths.
(3/3)(23/5) = 69/15
(5/5)(8/3) = 40/15
The problem is now the following:
69/15 - 40/15 = 29/15
Cheers,
Stan H.